# Digging into the Interior of Hot Cores with ALMA. VI. The Formation of Low-mass Multiple Systems in High-mass Cluster-forming Regions

### v1

@@ I Introduction
Multiple systems play a crucial role in stellar evolution, influencing processes from planet formation and orbital dynamics to supernova explosions and the formation of compact binaries, which are so far the most common sources of observed gravitational wave events (T. M. Tauris & E. P. J. van den Heuvel, 2023). The formation of multiple systems is generally explained by three pathways: core fragmentation, disk fragmentation, and dynamical capture (S. S. R. Offner et al., 2023). Core fragmentation occurs in the early stages, from the prestellar to the protostellar phase, where turbulence triggers multiple collapses individually within a single dense core on scales of thousands of au (S. P. Goodwin et al., 2004, 2007). Disk fragmentation takes place during the accretion phase, when parts of the accretion disk become gravitationally unstable (i.e., when the Toomre Q parameter 1) (J. E. Tohline, 2002; P. C. Clark et al., 2010; K. Kratter & G. Lodato, 2016). The resulting fragments typically form on scales of 100 au, consistent with disk sizes. Dynamical capture occurs when an initially unbound protostar becomes gravitationally bound to a nearby stellar system (E. C. Ostriker, 1994). Capturing the kinematic information of each protostar in multiple systems is challenging, so these formation pathways are usually distinguished by their characteristic fragmentation scales (S. S. R. Offner et al., 2023).
@@
Among the three, core fragmentation is considered a primary pathway for forming multiple systems (R. L. Kuruwita & T. Haugbølle, 2023). ALMA observations of fragmentation in starless cores have been conducted in nearby clouds such as Chamaeleon and Ophiuchus (M. M. Dunham et al., 2016; H. Kirk et al., 2017). S. Ohashi et al. (2018) and P. Caselli et al. (2019) identified multiple density peaks within individual prestellar cores, TUKH122 and L1544, respectively. More recently, D. Sahu et al. (2021) confirmed the presence of substructures within a prestellar core with sufficient mass ( 0.1 M⊙) and size ( 1200 au) to form protostars in the Orion cloud. Previous surveys toward the Perseus and Orion clouds have reported multiplicity fractions of 30% and 44% (J. J. Tobin et al., 2016, 2022) . These surveys reveal similar bimodal companion separation distributions, with one peak at 75 au, which likely results from disk fragmentation with a contribution of cores that formed via core fragmentation that migrated inward; while another peak at 4000 au is explained by primordial core fragmentation. Q.-y. Luo et al. (2022) suggests that both thermal Jeans fragmentation and turbulent fragmentation contribute to multiple system formation in star-forming cores in the Orion region predominantly forming low-mass stars. However, existing studies have mostly focused on nearby clouds (distance 500 pc) that only form low-mass stars.
@@
High-mass cluster-forming (HMCF) regions (1 pc structures forming both low- and high-mass stars) are considered representative sites of star formation in our Galaxy. Clump fragmentation from pc to a few thousand au in these regions is often explained by thermal Jeans fragmentation (P. Sanhueza et al., 2017, 2019; H. Beuther et al., 2018; K. Morii et al., 2023, 2024; K. Ishihara et al., 2024), however, how cores (a few 1000 au size objects) fragment into bound binary or higher-order multiple systems is unclear. The DIHCA (Digging into the Interior of Hot Cores with ALMA) survey covers 30 high-mass cluster-forming regions. S. Li et al. (2024) studied multiplicity in one examplar region, G333.23–0.06, and K. Taniguchi et al. (2023) identified hot cores in all 30 fields.
@@ II Observations
We observed 23 high-mass cluster-forming regions with ALMA between 2016/11/01 and 2018/11/25 in the C43-5 configuration, and between 2016/09/11 and 2019/11/15 in the C43-8 configuration, as part of the DIHCA project (F. A. Olguin et al., 2021, 2022, 2023, 2025; K. Taniguchi et al., 2023; S. Li et al., 2024; K. Ishihara et al., 2024; T. Sakai et al., 2025). The number of 12 m antennas used ranges from 41 to 46, sampling baselines up to 8500 m (Project IDs: 2016.1.01036.S, 2017.1.00237.S; PI: Patricio Sanhueza).
We used the Common Astronomy Software Application (CASA; (J. P. McMullin et al., 2007; CASA Team et al., 2022)) reduction pipeline to calibrate the data following the standard procedure. Since the data was obtained and assessed by the observatory in different cycles, each dataset was calibrated with its appropriate version of the CASA pipeline (versions 4.7.0, 4.7.2, 5.1.1-5, 5.4.0-70, and 5.6.1-8). The 1.3 mm continuum data is imaged using TCLEAN task of CASA. The line-free channels were identified by constructing dirty data cubes for each spectral window, extracting spectra at the averaged peak position, and applying asymmetric sigma clipping to remove channels affected by line emission, as detailed in (F. A. Olguin et al., 2021).
@@ III.1 Identification of Cores and Condensations
We have conducted high-angular resolution observations (005 and 03) of those cluster-forming regions in dust continuum emission at 1.33 mm using two ALMA configurations. We use mm wavelength observations because the large column of dust in high-mass cluster-forming regions precludes the direct detection of deeply embedded stellar embryos at shorter wavelengths (e.g., IR, optical, UV). At 1.33 mm and high resolution, we expect to detect the envelope and disk-like emission from young protostars. The compact configuration (03) captures spatial scales that allow us to define entities typically referred to as ‘cores’, that will form a single, a binary, or a higher-order multiple stellar system (K. Ishihara et al., 2024). In contrast, the extended configuration (005) provides the angular resolution to reveal the fragmentation within the core, tracing the cocoons that will ideally form single stars (referred here as ‘condensations’, J. E. Pineda et al., 2015).
@@
With an average spatial resolution of 1000 au, the compact ALMA configuration reveals a total of 415 cores that are very likely formed through thermal Jeans fragmentation (K. Ishihara et al., 2024). We therefore use the core properties obtained in Appendix A.1 and B.1 to separate the sample into high-mass and low-mass cores. We define a high-mass core as any core with mass higher than 16 M⊙ or cores that have characteristic ‘hot core’ emission lines, indicative of hosting a high-mass young stellar object. Assuming a star formation efficiency of 50% (C. D. Matzner & C. F. McKee, 2000), this corresponds to a threshold core mass of 16 M⊙, which we adopt as the dividing line between low- and high-mass cores. The identification of hot cores is based on K. Taniguchi et al. (2023), and these cores are not considered in this work. The subsequent analysis focuses on the cores that will likely form only low-mass stars (8 M⊙). With an average spatial resolution of 160 au, the newly presented high-resolution images reveal 446 condensations, either embedded within low-mass cores or located outside any identified core (See Appendix A.2).
@@ III.2 Identification of Multiple Systems
We combine two methods to define multiple systems. First, we search for multiple systems within cores, adopting the core boundary as the leaf structure decided by Astrodendro. The obtained core sizes are consistent with those from previous surveys of high-mass star-forming regions (F. Motte et al., 2022). Using the core boundary to search for companions/multiples will increase the likelihood of identifying gravitationally bound members. Adopting the core sizes normally observed in low-mass star-forming regions (0.1 pc) is unfeasible in the clustered environments observed in high-mass star-forming regions. This is because, in our sample, the large number of cores within 0.1 pc results in an increased likelihood of chance alignments, where unrelated condensations overlap along the line of sight and are mistakenly identified as multiple systems, as we have confirmed in one of the observed regions in S. Li et al. (2024). Second, for orphans (condensations without a parent core), we use the largest astrodendro leaf in each region to define an area as boundary to search for companions (details are described in Appendix D). Here, a companion is defined as any other condensation located within the same core, and within a defined area in the case of orphans. For the first time, we report the discovery of 72 low-mass multiple systems in high-mass cluster-forming regions.
@@ III.3 Multiplicity
Two informative metrics to characterize the multiplicity in star-forming regions are the multiplicity fraction (MF), which can be understood as the fraction of systems that are multiples, and the companion fraction (CF), understood as the fraction of companions per system (see details in Appendix E). We identified 52 multiple systems within the cores and 20 additional multiple systems among the orphan population. Including both cores and orphans, the overall distribution of systems is S:B:T:Q = 285:59:9:4. S, B, T, and Q represent single, binary, triple, and quadruple systems, respectively. For all the low-mass condensation population in high-mass cluster-forming regions, we obtain a MF and CF of 20% 2% and 25% 2% (Calculation seen in Appendix B). Condensations embedded in cores have a MF and a CF of 30% 3% and 38% 3%, respectively. Orphans are primarily single-systems with a MF and a CF of 10% 2% and 12% 2%, respectively. In the nearby, well-studied low-mass star-forming Perseus and Orion (including the northern Integral-Shaped Filament (ISF), L1641 and Orion B) clouds, previous surveys report a MF of about 30% with a resolution of 30 au (J. J. Tobin et al., 2022). To make a fair comparison with our survey, we recalculated their MF by treating systems with separations smaller than 160 au as a single system, corresponding to our average spatial resolution. The resulting MF in Perseus and Orion cloud is measured to be 17% 2% and 21% 2%, values that are consistent with those we obtain in high-mass cluster-forming regions.
@@
We adopt the method outlined by J. J. Tobin et al. (2022) to calculate the companion separation for multiple systems within cores, and to identify multiple systems in the orphan population (See Appendix F).
@@ IV.1 Physical Mechanisms of Core Fragmentation
Thermal Jeans fragmentation has been suggested to drive clump (1 pc) to core (few 1000s au) fragmentation in the high-mass cluster-forming regions under study (K. Ishihara et al., 2024). We examine whether it also plays a role in the fragmentation process leading to condensations (few 100s au). We assess this point following the work of Q.-y. Luo et al. (2022). They separate cores in Orion into two groups, those that fragment into multiples and those that do not. They show that the number density of the two groups are different, higher for cores that do fragment, in agreement with Jeans fragmentation (a higher density reduces the Jeans mass, promoting fragmentation). Following the same approach, Figure 3 compares the number density of cores forming single and multiple systems. The Kolmogorov-Smirnov (KS) test results show the p-value is 77, which indicates no significant difference between the number density distributions of the two groups. This result also implies that factors beyond thermal Jeans fragmentation, such as turbulence and/or magnetic fields (e.g., P. Sanhueza et al., 2025), play a more important role in the formation of low-mass multiple systems in high-mass cluster-forming regions.
@@ IV.2 Fragmentation Scale in High-mass Cluster-forming Regions
Surveys of nearby molecular clouds have revealed a bimodal companion separation distribution with one peak below 100 au and another around 4000 au (J. J. Tobin et al., 2022). These two peaks are widely interpreted as the imprint of distinct formation mechanisms, although this interpretation has been challenged by simulations (R. L. Kuruwita & T. Haugbølle, 2023). The closer peak (100 au) presumably arises from disk fragmentation caused by gravitational instabilities, while the larger peak reflects turbulent core fragmentation on scales of a few 1000s au within a collapsing core. The peak at 1200 au found in high-mass cluster-forming regions falls in between the two peaks observed in low-mass, nearby regions (Figure 2 (b)). Class 0 and Class I protostellar disks, as the kind of disks we would expect to find in our sample of low-mass condensations, have typical sizes smaller than 100 au (J. J. Tobin et al., 2020), unresolved at the resolution of our observations (G. Busquet et al., 2019). Furthermore, migration generally occurs inward, resulting in the contraction of member orbits (S. S. R. Offner et al., 2023; R. L. Kuruwita & T. Haugbølle, 2023). We can therefore discard disk fragmentation as the physical process producing a peak at 1200 au. This peak is likely equivalent to the 4000 au peak in low-mass regions produced by turbulent core fragmentation, but modified by the different environmental conditions at which high-mass stars form (Appendix I). We propose that the turbulent and denser environment of high-mass cluster-forming regions makes the most important difference.
@@
To investigate what defines the peak in the companion separation distribution, we first compare our observational results with numerical simulations. Self-gravitating radiation-hydrodynamics simulations (S. S. R. Offner et al., 2010; A. T. Lee et al., 2019) show that multiple systems formed via turbulent core fragmentation in clustered environments tend to have separations concentrated around 1000 au. Using magneto-hydrodynamics simulations, (R. L. Kuruwita & T. Haugbølle, 2023) explore how the initial separation of systems varies with mass/density of the hosting molecular cloud. These works find that the fragmentation scale appears to be smaller in denser environments and the peak of the initial separation distribution also has a maximum near 1000 au in the simulations with the highest density. Therefore, simulations provide further evidence that the peak of the separation distribution at 1200 au is produced by core fragmentation.
@@
To examine how the star-forming environment impacts the initial fragmentation separation, we assess the contribution of turbulence by defining a fragmentation scale via a Bonnor-Ebert analysis (following R. L. Kuruwita & T. Haugbølle, 2023). Rather than considering only thermal pressure and gravity (Jeans fragmentation), the Bonnor-Ebert radius of a core takes into account the pressure exerted by the turbulence and mass/density of the surrounding medium. Assuming that the typical radius at which a core fragments to produce multiples corresponding to this radius, we estimate a fragmentation scale within 600 300 au and 1000 500 au (See Appendix I), consistent with the observed peak of 1200 au from the observed companion separation distribution. We therefore propose that most of the low-mass multiple systems formed in these high-mass cluster-forming regions have been caught at an early moment in their evolution, near the moment of core fragmentation.
@@
Simulations suggest (S. S. R. Offner et al., 2010; A. T. Lee et al., 2019; R. L. Kuruwita & T. Haugbølle, 2023) that orbital migration can quickly occur in the order of 105 yr during the embedded stage. Migration inward is predicted to reduce the separation between multiple systems due to dynamical friction that dissipates energy and angular momentum (R. L. Kuruwita & T. Haugbølle, 2023). This process is particularly effective when the forming stars are not yet sufficiently massive with respect to the surrounding medium, and gas dynamics dominate. We interpret the asymmetry in the companion separation distribution, specifically the shallow slope toward smaller separations, as evidence for orbital migration (R. L. Kuruwita & T. Haugbølle, 2023). On the other hand, the steeper slope toward larger separations reflects the compactness of high-mass cluster formation (D. Guszejnov et al., 2023). The sharp drop in the distribution naturally arises from restricting the search of multiples to within the boundary of individual cores (or within 3000 au). If this limit is relaxed, in addition to increasing the likelihood of chance alignments, the association of multiples becomes contaminated by objects belonging to neighboring cores (See Appendix Figure 5(b)).

### v2

@@ I Introduction
Multiple systems play a crucial role in stellar evolution, influencing processes from planet formation and orbital dynamics to supernova explosions and the formation of compact binaries, which are so far the most common sources of observed gravitational wave events (Tauris and van den Heuvel, 2023). The formation of multiple systems is generally explained by three pathways: core fragmentation, disk fragmentation, and dynamical capture (Offner et al., 2023). Core fragmentation occurs in the early stages, from the prestellar to the protostellar phase, where turbulence triggers multiple collapses individually within a single dense core on scales of thousands of au (Goodwin et al., 2004, 2007). Disk fragmentation takes place during the accretion phase, when parts of the accretion disk become gravitationally unstable (i.e., when the Toomre Q parameter 1) (Tohline, 2002; Clark et al., 2010; Kratter and Lodato, 2016). The resulting fragments typically form on scales of 100 au, consistent with disk sizes. Dynamical capture occurs when an initially unbound protostar becomes gravitationally bound to a nearby stellar system (Ostriker, 1994). Capturing the kinematic information of each protostar in multiple systems is challenging, so these formation pathways are usually distinguished by their characteristic fragmentation scales (Offner et al., 2023).
@@
Among the three, core fragmentation is considered a primary pathway for forming multiple systems (Kuruwita and Haugbølle, 2023). ALMA observations of fragmentation in starless cores have been conducted in nearby clouds such as Chamaeleon and Ophiuchus (Dunham et al., 2016; Kirk et al., 2017). Ohashi et al. (2018) and Caselli et al. (2019) identified multiple density peaks within individual prestellar cores, TUKH122 and L1544, respectively. More recently, Sahu et al. (2021) confirmed the presence of substructures within a prestellar core with sufficient mass ( 0.1 M⊙) and size ( 1200 au) to form protostars in the Orion cloud. Previous surveys toward the Perseus and Orion clouds have reported multiplicity fractions of 30% and 44% (Tobin et al., 2016, 2022) . These surveys reveal similar bimodal companion separation distributions, with one peak at 75 au, which likely results from disk fragmentation with a contribution of cores that formed via core fragmentation that migrated inward; while another peak at 4000 au is explained by primordial core fragmentation. Luo et al. (2022) suggests that both thermal Jeans fragmentation and turbulent fragmentation contribute to multiple system formation in star-forming cores in the Orion region predominantly forming low-mass stars. However, existing studies have mostly focused on nearby clouds (distance 500 pc) that only form low-mass stars.
@@
High-mass cluster-forming (HMCF) regions (1 pc structures forming both low- and high-mass stars) are considered representative sites of star formation in our Galaxy. Clump fragmentation from pc to a few thousand au in these regions is often explained by thermal Jeans fragmentation (Sanhueza et al., 2017, 2019; Beuther et al., 2018; Morii et al., 2023, 2024; Ishihara et al., 2024), however, how cores (a few 1000 au size objects) fragment into bound binary or higher-order multiple systems is unclear. The DIHCA (Digging into the Interior of Hot Cores with ALMA) survey covers 30 high-mass cluster-forming regions. Li et al. (2024) studied multiplicity in one examplar region, G333.23–0.06, and Taniguchi et al. (2023) identified hot cores in all 30 fields.
@@ II Observations
We observed 23 high-mass cluster-forming regions with ALMA between 2016/11/01 and 2018/11/25 in the C43-5 configuration, and between 2016/09/11 and 2019/11/15 in the C43-8 configuration, as part of the DIHCA project (Olguin et al., 2021, 2022, 2023, 2025; Taniguchi et al., 2023; Li et al., 2024; Ishihara et al., 2024; Sakai et al., 2025). The number of 12 m antennas used ranges from 41 to 46, sampling baselines up to 8500 m (Project IDs: 2016.1.01036.S, 2017.1.00237.S; PI: Patricio Sanhueza).
We used the Common Astronomy Software Application (CASA; (McMullin et al., 2007; CASA Team et al., 2022)) reduction pipeline to calibrate the data following the standard procedure. Since the data was obtained and assessed by the observatory in different cycles, each dataset was calibrated with its appropriate version of the CASA pipeline (versions 4.7.0, 4.7.2, 5.1.1-5, 5.4.0-70, and 5.6.1-8). The 1.3 mm continuum data is imaged using TCLEAN task of CASA. The line-free channels were identified by constructing dirty data cubes for each spectral window, extracting spectra at the averaged peak position, and applying asymmetric sigma clipping to remove channels affected by line emission, as detailed in (Olguin et al., 2021).
@@ III.1 Identification of Cores and Condensations
We have conducted high-angular resolution observations (005 and 03) of those cluster-forming regions in dust continuum emission at 1.33 mm using two ALMA configurations. We use mm wavelength observations because the large column of dust in high-mass cluster-forming regions precludes the direct detection of deeply embedded stellar embryos at shorter wavelengths (e.g., IR, optical, UV). At 1.33 mm and high resolution, we expect to detect the envelope and disk-like emission from young protostars. The compact configuration (03) captures spatial scales that allow us to define entities typically referred to as ‘cores’, that will form a single, a binary, or a higher-order multiple stellar system (Ishihara et al., 2024). In contrast, the extended configuration (005) provides the angular resolution to reveal the fragmentation within the core, tracing the cocoons that will ideally form single stars (referred here as ‘condensations’, Pineda et al., 2015).
@@
With an average spatial resolution of 1000 au, the compact ALMA configuration reveals a total of 415 cores that are very likely formed through thermal Jeans fragmentation (Ishihara et al., 2024). We therefore use the core properties obtained in Appendix A.1 and B.1 to separate the sample into high-mass and low-mass cores. We define a high-mass core as any core with mass higher than 16 M⊙ or cores that have characteristic ‘hot core’ emission lines, indicative of hosting a high-mass young stellar object. Assuming a star formation efficiency of 50% (Matzner and McKee, 2000), this corresponds to a threshold core mass of 16 M⊙, which we adopt as the dividing line between low- and high-mass cores. The identification of hot cores is based on Taniguchi et al. (2023), and these cores are not considered in this work. The subsequent analysis focuses on the cores that will likely form only low-mass stars (8 M⊙). With an average spatial resolution of 160 au, the newly presented high-resolution images reveal 446 condensations, either embedded within low-mass cores or located outside any identified core (See Appendix A.2).
@@ III.2 Identification of Multiple Systems
We combine two methods to define multiple systems. First, we search for multiple systems within cores, adopting the core boundary as the leaf structure decided by Astrodendro. The obtained core sizes are consistent with those from previous surveys of high-mass star-forming regions (Motte et al., 2022). Using the core boundary to search for companions/multiples will increase the likelihood of identifying gravitationally bound members. Adopting the core sizes normally observed in low-mass star-forming regions (0.1 pc) is unfeasible in the clustered environments observed in high-mass star-forming regions. This is because, in our sample, the large number of cores within 0.1 pc results in an increased likelihood of chance alignments, where unrelated condensations overlap along the line of sight and are mistakenly identified as multiple systems, as we have confirmed in one of the observed regions in Li et al. (2024). Second, for orphans (condensations without a parent core), we use the largest astrodendro leaf in each region to define an area as boundary to search for companions (details are described in Appendix D). Here, a companion is defined as any other condensation located within the same core, and within a defined area in the case of orphans. For the first time, we report the discovery of 72 low-mass multiple systems in high-mass cluster-forming regions.
@@ III.3 Multiplicity
Two informative metrics to characterize the multiplicity in star-forming regions are the multiplicity fraction (MF), which can be understood as the fraction of systems that are multiples, and the companion fraction (CF), understood as the fraction of companions per system (see details in Appendix E). We identified 52 multiple systems within the cores and 20 additional multiple systems among the orphan population. Including both cores and orphans, the overall distribution of systems is S:B:T:Q = 285:59:9:4. S, B, T, and Q represent single, binary, triple, and quadruple systems, respectively. For all the low-mass condensation population in high-mass cluster-forming regions, we obtain a MF and CF of 20% 2% and 25% 2% (Calculation seen in Appendix B). Condensations embedded in cores have a MF and a CF of 30% 3% and 38% 3%, respectively. Orphans are primarily single-systems with a MF and a CF of 10% 2% and 12% 2%, respectively. In the nearby, well-studied low-mass star-forming Perseus and Orion (including the northern Integral-Shaped Filament (ISF), L1641 and Orion B) clouds, previous surveys report a MF of about 30% with a resolution of 30 au (Tobin et al., 2022). To make a fair comparison with our survey, we recalculated their MF by treating systems with separations smaller than 160 au as a single system, corresponding to our average spatial resolution. The resulting MF in Perseus and Orion cloud is measured to be 17% 2% and 21% 2%, values that are consistent with those we obtain in high-mass cluster-forming regions.
@@
We adopt the method outlined by Tobin et al. (2022) to calculate the companion separation for multiple systems within cores, and to identify multiple systems in the orphan population (See Appendix F).
@@ IV.1 Physical Mechanisms of Core Fragmentation
Thermal Jeans fragmentation has been suggested to drive clump (1 pc) to core (few 1000s au) fragmentation in the high-mass cluster-forming regions under study (Ishihara et al., 2024). We examine whether it also plays a role in the fragmentation process leading to condensations (few 100s au). We assess this point following the work of Luo et al. (2022). They separate cores in Orion into two groups, those that fragment into multiples and those that do not. They show that the number density of the two groups are different, higher for cores that do fragment, in agreement with Jeans fragmentation (a higher density reduces the Jeans mass, promoting fragmentation). Following the same approach, Figure 3 compares the number density of cores forming single and multiple systems. The Kolmogorov-Smirnov (KS) test results show the p-value is 77, which indicates no significant difference between the number density distributions of the two groups. This result also implies that factors beyond thermal Jeans fragmentation, such as turbulence and/or magnetic fields (e.g., Sanhueza et al., 2025), play a more important role in the formation of low-mass multiple systems in high-mass cluster-forming regions.
@@ IV.2 Fragmentation Scale in High-mass Cluster-forming Regions
Surveys of nearby molecular clouds have revealed a bimodal companion separation distribution with one peak below 100 au and another around 4000 au (Tobin et al., 2022). These two peaks are widely interpreted as the imprint of distinct formation mechanisms, although this interpretation has been challenged by simulations (Kuruwita and Haugbølle, 2023). The closer peak (100 au) presumably arises from disk fragmentation caused by gravitational instabilities, while the larger peak reflects turbulent core fragmentation on scales of a few 1000s au within a collapsing core. The peak at 1200 au found in high-mass cluster-forming regions falls in between the two peaks observed in low-mass, nearby regions (Figure 2 (b)). Class 0 and Class I protostellar disks, as the kind of disks we would expect to find in our sample of low-mass condensations, have typical sizes smaller than 100 au (Tobin et al., 2020), unresolved at the resolution of our observations (Busquet et al., 2019). Furthermore, migration generally occurs inward, resulting in the contraction of member orbits (Offner et al., 2023; Kuruwita and Haugbølle, 2023). We can therefore discard disk fragmentation as the physical process producing a peak at 1200 au. This peak is likely equivalent to the 4000 au peak in low-mass regions produced by turbulent core fragmentation, but modified by the different environmental conditions at which high-mass stars form (Appendix I). We propose that the turbulent and denser environment of high-mass cluster-forming regions makes the most important difference.
@@
To investigate what defines the peak in the companion separation distribution, we first compare our observational results with numerical simulations. Self-gravitating radiation-hydrodynamics simulations (Offner et al., 2010; Lee et al., 2019) show that multiple systems formed via turbulent core fragmentation in clustered environments tend to have separations concentrated around 1000 au. Using magneto-hydrodynamics simulations, (Kuruwita and Haugbølle, 2023) explore how the initial separation of systems varies with mass/density of the hosting molecular cloud. These works find that the fragmentation scale appears to be smaller in denser environments and the peak of the initial separation distribution also has a maximum near 1000 au in the simulations with the highest density. Therefore, simulations provide further evidence that the peak of the separation distribution at 1200 au is produced by core fragmentation.
@@
To examine how the star-forming environment impacts the initial fragmentation separation, we assess the contribution of turbulence by defining a fragmentation scale via a Bonnor-Ebert analysis (following Kuruwita and Haugbølle, 2023). Rather than considering only thermal pressure and gravity (Jeans fragmentation), the Bonnor-Ebert radius of a core takes into account the pressure exerted by the turbulence and mass/density of the surrounding medium. Assuming that the typical radius at which a core fragments to produce multiples corresponding to this radius, we estimate a fragmentation scale within 600 300 au and 1000 500 au (See Appendix I), consistent with the observed peak of 1200 au from the observed companion separation distribution. We therefore propose that most of the low-mass multiple systems formed in these high-mass cluster-forming regions have been caught at an early moment in their evolution, near the moment of core fragmentation.
@@
Simulations suggest (Offner et al., 2010; Lee et al., 2019; Kuruwita and Haugbølle, 2023) that orbital migration can quickly occur in the order of 105 yr during the embedded stage. Migration inward is predicted to reduce the separation between multiple systems due to dynamical friction that dissipates energy and angular momentum (Kuruwita and Haugbølle, 2023). This process is particularly effective when the forming stars are not yet sufficiently massive with respect to the surrounding medium, and gas dynamics dominate. We interpret the asymmetry in the companion separation distribution, specifically the shallow slope toward smaller separations, as evidence for orbital migration (Kuruwita and Haugbølle, 2023). On the other hand, the steeper slope toward larger separations reflects the compactness of high-mass cluster formation (Guszejnov et al., 2023). The sharp drop in the distribution naturally arises from restricting the search of multiples to within the boundary of individual cores (or within 3000 au). If this limit is relaxed, in addition to increasing the likelihood of chance alignments, the association of multiples becomes contaminated by objects belonging to neighboring cores (See Appendix Figure 5(b)).